OpenStudy (anonymous):

A cylindrical tank with diameter 30 ft is filled with gasoline to a depth of 60ft. The gasoline begins draining at a constant rate of 5 cubic feet per sec. Write the volume of gasoline remaining in the tank t seconds after the tank begins draining as a function of t.

OpenStudy (anonymous):

Okay, so we start with a volume of gasoline that can be represented as the volume of a cylinder with diameter 30ft and height 60ft. If diameter is 30ft then the area of the base is 225\(\pi\) square feet, multiply by the height to get total volume of \(13,500\pi\) cubic feet. Then it's draining at 5 cubic feet per second, so the final expression is \(13,500\pi-5t\) cubic feet. Yeah?

OpenStudy (anonymous):

I think that would be right